OpenStudy (anonymous):

Solve: Question 1: 3(5)^2x-3=6 Question 2: 2^x-4+10=22 Question 3: log_3 3+log_3 x=5.

OpenStudy (campbell_st):

so in the 1st question divide both sides of the equation by 3 and you get \[5^{2x - 3} = 2\] now take the log of both sides... base e or base 10 \[\ln(5^{2x -3}) = \ln(2)\] apply the log law for powers \[(2x -3)\ln(5) = \ln(2)\] now you should be able to solve for x.